The expression 52 - 4 of (17 - 12) + 4 × 7 is equal to


1
Expert's answer
2021-05-07T10:08:23-0400

"\text{ The expression }52 - 4 \text{ of } (17 - 12) + 4 × 7 \text{ is}"


"= 52-4\times (5)+28

 =52-20+28=60"